I was given this problem. I am in prep math classes so trying to get back in trhe swing of things.
y= 1/1-x
x=.8
solve for y

Answer by LinnW(1048)

Replace .8 for x in y=1/(1-x)
y=1/(1-.8)
y=1/(1.0-0.8)
y=1/(0.2)
y=5